OpenStudy (anonymous):

What is the local maximum value of the function? (Round answer to the nearest thousandth.) g(x) = 3x^3 + 3x^2 - 30x + 24

OpenStudy (superdavesuper):

u need to know differentiation to solve this: at local max, g'(x)=0. solve for x then u will find the value of g(x) there.

OpenStudy (mathmale):

supernerd: Please find the derivative of the given function g(x) and share your result here.

OpenStudy (anonymous):

For the derivative I got: 9x^2+6x-30 @mathmale

OpenStudy (mathmale):

Great. Now, as Dave has already suggested, please set that = to 0 and solve for x. You may find it easier to factor out the common factor 3 before you do this.

OpenStudy (mathmale):

Hint: the results of using the quadratic formula in this case to find roots are not pretty!

OpenStudy (superdavesuper):

thanks @mathmale @supernerd it may not be pretty but g'(x)=0 is solvable; here is how g(x) look like :)

OpenStudy (anonymous):

g(x)=

OpenStudy (anonymous):

g'(x)=9x^2+6x-30 9x^2+6x-30=0 critical vaule= x=-(sqrt(31)+1)/3 or x=(sqrt(31)-1)/3

OpenStudy (anonymous):

I simplified those all further and got 1.5226, is that correct?

OpenStudy (mathmale):

How could you check this result for yourself? Doing so would make you a supernerd. Hint: Graph the given function and then vary x until you get x= 1.523. See a max or a min? Of course there are other ways to do this; you could use synth div, but that would be messy.

OpenStudy (superdavesuper):

@mathmale is right if u look at the attached graph to my previous post, u can see that the max is quite different. so i think u made a mistake somewhere.

OpenStudy (superdavesuper):

oh wait @supernerd i think u found the local min instead of the max...try the other root for g'(x)=0 plz

OpenStudy (mathmale):

Right, Dave; I did a quick calculator calculation of the value of y at supernerd's x=1.523 and found that the resulting point, (1.523,-4.133), is indeed a relative minimum of the given function.

OpenStudy (superdavesuper):

hehehe i cheated....i used the web site that begins w "w" n thats where i got the graph too ;)

OpenStudy (mathmale):

I think that's a fine approach; we know how to graph manually, so if we can do it much more quickly online, why not?

OpenStudy (anonymous):

Ah, I keep trying but I just keep getting the same answer, how do we find the maximum as opposed to the minimum?

OpenStudy (superdavesuper):

there should be two roots to g'(x)=0 as it is a quadratic eqn. 1 is the max n the other is the min. So try the other root plz :)

OpenStudy (anonymous):

I got 2.1893! Is that correct? @superdavesuper

OpenStudy (mathmale):

Not to be unkind, but I think you should be thinking about how y ou could check your own answers. See Dave's illustration: http://assets.openstudy.com/updates/attachments/52e404f8e4b096fa40598f39-superdavesuper-1390677131242-untitled.png and ask yourself whether or not this graph seems to have a local max when x=2.1893.

OpenStudy (mathmale):

"critical value= x=-(sqrt(31)+1)/3 or x=(sqrt(31)-1)/3" came from "annasjuice." Try evaluating these again. One value should be 1.523; the other one should be (-).

OpenStudy (anonymous):

Oh! This time I got: -2.1893

OpenStudy (mathmale):

A lot better! Now how are you going to determine whether or not that's "right"?

OpenStudy (anonymous):

plug it in to the original equation? Or graph it?

OpenStudy (mathmale):

Look at Dave's original illustration. Does it appear that his graph has a max at your -2.189? http://assets.openstudy.com/updates/attachments/52e404f8e4b096fa40598f39-superdavesuper-1390677131242-untitled.png

OpenStudy (anonymous):

Yes!